An oxide m2o3 has 68% metal, then what is the atomic weight of metal

Grade:11

1 Answers

Arun
5 years ago
Dear Shruti
 
let us assume
atomic weight = m
 
now 0.68 = 2m / (2m + 48)
 
1.36 m + 32.64 = 2m
 
0.64 m = 32.64
 
m = 32.64 / 0.64 = 51
 
hence atomic weight of metal  = 51
 
Regards
Arun (askIITians forum expert)